solving quadratic equations by completing the square

HELP PLEASE

x^2+4+1=0
or like
x^2+x-1=0

thank you!!

Answer by nerdybill(7384)

You forgot the x-term:
I'm assuming instead of:
x^2+4+1=0
you meant:
x^2+4x+1=0
x^2+4x=-1
half b then double -- add to both sides:
x^2+4x+4=-1+4
(x+2)^2 = 3
x+2 = sqrt(3)
x = +-sqrt(3) - 2
x = {-3.732, -0.268}
.
.
x^2+x-1=0
x^2+x = 1
x^2+x+ 1/4 = 1 + 1/4
(x+ 1/2)^2 = 1.25
(x+.5)^2 = 1.25
x+.5 = sqrt(1.25)
x = +-sqrt(1.25) - .5
x = {-1.618, 0.618}